# Bootstrapping the 3d Ising Stress Tensor

### v1

@@ 1 Introduction
However, there are several reasons why we would like to go beyond these results. Firstly, it is not clear that the allowed region of Kos:2016ysd can be made systematically smaller by increasing the derivative order of bootstrap functionals. Relatedly, there are certain low-lying Regge trajectories of operators that are expected theoretically Fitzpatrick:2012yx ; Komargodski:2012ek ; Fitzpatrick:2015qma but not present in the extremal spectra studied in Simmons-Duffin:2016wlq . These include the $[\sigma T]$, $[\epsilon T]$, and $[TT]$ double-twist trajectories built out of the stress-energy tensor $T^{\mu\nu}$ (which can also be interpreted as triple-twist or quadruple-twist trajectories built out of $\sigma$). Furthermore, the spectra of Simmons-Duffin:2016wlq give no information about $\mathbb{Z}_{2}$-even operators of odd spin, parity-odd operators, or operator product expansion (OPE) coefficients involving multiple spinning operators (e.g. those appearing in $\langle TT\epsilon\rangle$ or $\langle TTT\rangle$).
@@
Table 1: Bounds on scaling dimensions and OPE coefficients in the 3d Ising CFT, using the numerical bootstrap for correlators of $\sigma$, $\epsilon$, and the stress tensor $T$. Here, $c_{T}$ (the “central charge”) is the coefficient of the stress-tensor two-point function, and $n_{B}$ and $n_{F}$ are OPE coefficients in the stress-tensor three-point function. The OPE coefficient $\lambda_{TT\epsilon}$ is defined in section 2.1. Our bounds are computed using derivative order $\Lambda=43$. Bold-face errors are rigorous. We also include comparisons to previous bootstrap calculations and the best determinations we are aware of from other numerical methods. Note: the errors in Kos:2016ysd have been slightly underestimated, see footnote 2. ∗∗Note: the results for $(n_{B},n_{F})$ in Dymarsky:2017yzx assume a parity-odd scalar gap $\Delta_{-}\geq 10$. The weaker assumption $\Delta_{-}\geq 3$ gave $(n_{B},n_{F})=(0.919(18),0.027(18))$.
@@
Figure 1: Allowed regions in $(\Delta_{\sigma},\Delta_{\epsilon})$-space determined from the $T\sigma\epsilon$ system of correlators at $\Lambda=27,35,43$ (orange, green, and red), together with the allowed region determined from the $\sigma\epsilon$ system at $\Lambda=43$ (blue) Kos:2016ysd . [^2] The $\sigma\epsilon$ allowed region shown in this figure was re-computed for this work and is a more accurate version of the allowed region from Kos:2016ysd . Compared to Kos:2016ysd , we have benefitted from the OPE cutting surface algorithm Chester:2019ifh , improvements in computational efficiency, and more powerful computational resources. Our $\sigma\epsilon$ allowed region is slightly larger than the one in Kos:2016ysd but this does not significantly affect any of their results. We thank Slava Rychkov and Ning Su for discussions that motivated us to perform this calculation. On the left, we show a wider view where all the regions are visible. On the right, we zoom in on the $T\sigma\epsilon$ regions at $\Lambda=35$ and $\Lambda=43$. At $\Lambda=43$, the $T\sigma\epsilon$ allowed region is roughly $40$-50×50$\times$50 × smaller than the $\sigma\epsilon$ allowed region.
A natural continuation of these works is to consider a mixed-correlator system involving all nonvanishing 4-point functions of $\sigma$, $\epsilon$, and $T^{\mu\nu}$. While such a large system comes with technical challenges, they are not insurmountable and we will see that combining these constraints allows us to make precise predictions for CFT data involving the stress tensor, opening up a new era of precision in the numerical study of the 3d Ising CFT. Along the way we have significantly improved several software tools and algorithms used for the numerical bootstrap. For the impatient reader, a concise summary of our new results for scaling dimensions and OPE coefficients is provided in table 1. We also provide a plot of the allowed island in $(\Delta_{\sigma},\Delta_{\epsilon})$-space (which is about 50 times smaller than the previous state of the art) in figure 1.
@@ 2.1 Three-point structures
[…]
@@
We emphasize that it is important to ensure that three-point structures for a given channel are linearly-independent for all allowed values of $\Delta$ [^4] We observed empirically that one only needs to ensure linear independence for the values of $\Delta$ allowed by unitarity and the gap assumptions. Presumably this happens because in SDPB the conformal blocks are eventually sampled at a large number of allowed values of $\Delta$ and polynomially interpolated. This interpolation will very accurately reproduce any degeneracies in the allowed region of $\Delta$, but will likely break any degeneracies outside of this region where the polynomial interpolation is of poor quality. Without care, it can happen that the space spanned by a naïve choice of conserved structures is degenerate at specific values of $\Delta$. This occurs, for example, if one solves the conservation equations in Mathematica in a generic basis.
@@ 2.2 Ward identities
[…]
@@ 2.3 Four-point structures and crossing equations
To describe four-point tensor structures we use the $q$-basis of Kravchuk:2016qvl , as it does not suffer from redundancies which plague embedding-space bases Costa:2011mg ; Dymarsky:2013wla . A $q$-basis four-point tensor structure is denoted by $[q_{1}q_{2}q_{3}q_{4}]^{\pm}$, where $q_{i}\in\{-\ell_{i},-\ell_{i}+1,\dots,+\ell_{i}\}$ and $\ell_{i}$ is the spin of the operator at position $i$. The sign $\pm$ indicates the charge of the tensor structure under exchange of the cross-ratios $z\leftrightarrow\overline{z}$.
@@
$\Pi_{\langle TTTT\rangle}$ =$=\{\text{id},(12)(34),(13)(24),(14)(23)\},$ (16)
@@
with analogous expressions for other orderings. Any $q$-basis structure can be symmetrized under $\Pi$, and we denote symmetrization of $[q_{1}q_{2}q_{3}q_{4}]^{\pm}$ by $\langle q_{1}q_{2}q_{3}q_{4}\rangle^{\pm}$. Four-point functions can now be written as
@@ 3.2 Optimizing blocks_3d
In order to explain these improvements, we must give a brief overview of the function of blocks_3d. A single run of blocks_3d computes multiple conformal blocks, which all share the same external operator representations, four-point structure label, as well as the $j_{12}$ and $j_{43}$ parts of the three-point function labels. The program outputs conformal blocks with all possible values of $j_{120}$ and $j_{430}$, where the index 00 refers to the exchanged operator, as well as all values of the exchanged spin $j$ up to a cutoff $j_{\text{max}}$.
@@
These optimizations were implemented over a period of time of continuous development, and we were using the software the entire time. As a result, we have not performed a systematic comparison of the original version to the final version used in this project. However, in comparisons between intermediate versions we observed, for reasonably representative runs, a more than 10$\times$ improvement in memory usage and a similar improvement in computation (wall) time. Note that the improvements in memory usage also mean that more blocks_3d instances can run on a single machine. This is parallelism with nearly perfect scaling, unlike the parallelization within a single blocks_3d instance, [^14] This does not contradict the previously stated importance of this internal parallelization. further reducing the time needed to generate all conformal blocks.
@@ 4.1 Review: cutting surface algorithm
Importantly, this constraint is unchanged if we rescale $\vec{\lambda}$ by a non-zero real number. Therefore, it only depends on the equivalence class $[\vec{\lambda}]$ of $\vec{\lambda}$ in $\mathbb{R}\mathbb{P}^{4}$ and we should treat $[\vec{\lambda}]$ as the new parameter. If we are able to exclude all possible values of $[\vec{\lambda}]$ for a fixed $(\Delta_{\sigma},\Delta_{\epsilon})$, then the point $(\Delta_{\sigma},\Delta_{\epsilon})$ is excluded in the dimension space. As shown in Kos:2016ysd , this procedure yields a significantly smaller allowed region for $(\Delta_{\sigma},\Delta_{\epsilon})$ than using (31).
How can one exclude all possible values of $[\vec{\lambda}]$ in a finite amount of time? The approach used in Kos:2016ysd was to discretize $\mathbb{R}\mathbb{P}^{1}$ (their $\vec{\lambda}$ had only two components) using a finite grid. If all points in the grid are excluded, one can be reasonably confident that there are no allowed points in $\mathbb{R}\mathbb{P}^{1}$ if the grid is fine enough. This approach worked well in Kos:2013tga because their parameter space $\mathbb{R}\mathbb{P}^{1}$ was one-dimensional. In our case the parameter space is $\mathbb{R}\mathbb{P}^{4}$ and discretization appears computationally infeasible.
This problem was solved in Chester:2019ifh with the “cutting surface” algorithm. The key observation is simply that a finite set of functionals $\alpha_{I}^{(k)}$, $k=\{1,2,\dots,N\}$, can be enough to exclude the full parameter space $\mathbb{R}\mathbb{P}^{4}$. Indeed, existence of a functional $\alpha_{I}^{(k)}$ implies that all $[\vec{\lambda}]$ which satisfy
@@
If (36) holds, then all $[\vec{\lambda}]$ are excluded. If (36) doesn’t hold, then we can take a $[\vec{\lambda}_{N+1}]$ that is not in $\bigcup_{k=1}^{N}E_{k}$, i.e.
@@
If such a functional $\alpha_{I}^{(N+1)}$ is found, we can add it to the list, increment $N$, and check (36) again. If a functional cannot be found, it means that $[\vec{\lambda}_{N+1}]$ is allowed and so are the current values of $(\Delta_{\sigma},\Delta_{\epsilon})$. Typically, this process converges in a number of steps which is much smaller than the number of points in a fine enough grid on $\mathbb{R}\mathbb{P}^{4}$. It also has the benefit of not relying on a discretization.
The technical content of the cutting surface algorithm Chester:2019ifh is in the procedure for finding a solution $[\vec{\lambda}_{N+1}]$ to (37) or determining that no solution exists. [^17] When several solutions to (37) can be found, the cutting surface algorithm will try to select a solution that is the furthest away from the sets $E_{k}$ in order to speed up the convergence of the whole process. Abstractly, this can be formulated as the following problem: given quadratic forms $Q_{0},Q_{1},\dots,Q_{N}$, determine whether there exists a solution $\vec{\lambda}\in\mathbb{R}^{n+1}$ to the equations
@@
the quadratic form $Q_{0}$ is specified manually to restrict the search to an ellipse in $\mathbb{R}\mathbb{P}^{n+1}$. This bounding ellipse is chosen conservatively to include the region where the allowed points $[\vec{\lambda}]$ are expected to be. See section 5.2 for details. In essence, one of our assumptions is that if any $[\vec{\lambda}]$ is allowed at all, it is deep inside the region defined by $\vec{\lambda}^{T}Q_{0}\vec{\lambda}<0$. The technical meaning of this will be clarified below.
@@ 4.2 Certification algorithm
Out of the 38 disallowed points in $(\Delta_{\sigma},\Delta_{\epsilon})$ space that we computed at $\Lambda=43$, 30 points were certified by $\mathcal{B}[\delta c^{1}]=\emptyset$, 1 point by $\mathcal{B}[\delta c^{2}]=\emptyset$, and 2 points by by $\mathcal{B}[\delta c^{3}]=\emptyset$. This produces certificates for 33 out of 38 points.
@@ 5.1.1 Bounds on scalar gaps
Figure 4: Bound on the allowed gaps for parity-even and -odd scalars in the single-correlator $\langle TTTT\rangle$ system. Successively interior regions correspond to higher derivative order $\Lambda$. We can see that by $\Lambda=43$ the bounds are fairly well converged and therefore are hard to distinguish visually at this scale. The known values for these gaps in the free scalar (FS) theory, Majorana free fermion (FF) theory, stress tensor mean-field theory (MFT), $O(N)$ CFT in the large-$N$ limit ($O(\infty)$), and Ising CFT are indicated.
@@ 5.2 Mixed correlators of $T,\sigma,\epsilon$
$+$ $+$ 0 $\Delta\geq 3$ $\Delta=3.82951(61)$ Reehorst:2021hmp
-- $+$ 0 $\Delta\geq 3$ $\Delta=5.262(89)$ Reehorst:2021hmp
$+$ -- 0 $\Delta\geq 3$ $\Delta\gtrsim 10$ Zhu:2022gjc
$+$ $+$ 2 $\Delta\geq 4$ $\Delta=5.499(17)$ Reehorst:2021hmp
@@
Table 6: Our gap assumptions for operators other than $\sigma,\epsilon,T$, used for computing the allowed island in the 3d Ising CFT. Essentially, we assume that all scalars (other than $\sigma$, $\epsilon$) are irrelevant, and that there is a gap of $1$ above the stress tensor. We also assume a small twist gap in all other sectors ($\tau=\Delta-\ell$). In the right-hand column, we display the expected actual values for these gaps in the 3d Ising CFT. In the last row of the right-hand column, the twist gap comes from the $\mathbb{Z}_{2}$-even, parity-even, spin-4 sector, which has $\Delta_{C_{\mu\nu\rho\sigma}}=5.022665(28)$. Clearly our gap assumptions are very conservative.
At lower derivative orders we computed bounds following the navigator and skydiving approach described above. We found the minimum of the navigator $\mathcal{N}_{\min}$ and maximum and minimum values of $\Delta_{\sigma}$ across the space of external data. For example, at $\Lambda=11,19$ we obtained the results given below.
@@
To compute bounds at higher derivative order, we followed a similar procedure to the computations of bootstrap islands in the $O(2)$ model Chester:2019ifh , $O(3)$ model Chester:2020iyt , and GNY models Erramilli:2022kgp . For each point in $(\Delta_{\sigma},\Delta_{\epsilon})$-space, we use the OPE “cutting surface” algorithm of Chester:2019ifh to find an allowed direction for the vector $\vec{\lambda}$. The output of this algorithm is either an equivalence class $[\vec{\lambda}]\in\mathbb{RP}^{4}$ for which the corresponding SDP is primal feasible, or evidence that no such $[\vec{\lambda}]$ exists (see section 4 for details). In the former case, we say that the point $(\Delta_{\sigma},\Delta_{\epsilon})$ is “allowed,” and in the latter case we say it is “disallowed.”
@@
[…]
@@
The OPE cutting surface algorithm requires an initial bounding ellipse to work efficiently. We determine it by working our way up in derivative order $\Lambda$, with a method similar to our strategy for finding affine coordinates $(x,y)$. At a given value of $\Lambda$, we collect together all the allowed points in OPE space $[\vec{\lambda}]$ and use principal component analysis (in an affine patch of $\mathbb{RP}^{4}$) to find an ellipse that contains them all. Then we expand this ellipse by an $O(1)$ factor, say $4$. Finally, we use the expanded ellipse as the bounding ellipse for the next higher value $\Lambda^{\prime}>\Lambda$. As a consistency check, we observe afterwards that the allowed OPE points $[\vec{\lambda}]$ at the higher $\Lambda^{\prime}$ cluster in a small region deep in the interior of the bounding ellipse, indicating that the results are independent of the choice of bounding ellipse. As an example, the bounding ellipse at $\Lambda=43$ is defined by $\vec{\lambda}^{T}Q_{0}\vec{\lambda}\leq 0$, where $Q_{0}$ is given by the matrix
@@
In figure 8, we show a plot of the allowed points $[\vec{\lambda}]$ at $\Lambda=43$, compared to this bounding ellipse (which is a sphere of radius $4$ in the coordinates of the plot). As claimed, all the allowed points are deep in the interior.
Figure 8: Allowed values of $[\vec{\lambda}]$ found at $\Lambda=43$, compared to the initial bounding ellipse, in coordinates where the bounding ellipse is a sphere of radius $4$. Here, we show just 3 out of 4 coordinates of $[\vec{\lambda}]$ — the other coordinate is similar. All the allowed points are deep inside the bounding ellipse.
As mentioned above, each allowed point in $(\Delta_{\sigma},\Delta_{\epsilon})$-space comes with an equivalence class $[\vec{\lambda}]\in\mathbb{RP}^{4}$, determined by the OPE cutting surface algorithm. Next, for each allowed point, we compute upper and lower bounds on the norm-squared $|\vec{\lambda}|^{2}$ using the standard procedure for bounding OPE coefficients Caracciolo:2009bx . This results in pairs of OPE coefficient vectors $\vec{\lambda}_{u},\vec{\lambda}_{l}$ with norms given by the upper/lower bounds on $|\vec{\lambda}|$. Collecting together the data $(\Delta_{\sigma},\Delta_{\epsilon},\vec{\lambda}_{u})$ and $(\Delta_{\sigma},\Delta_{\epsilon},\vec{\lambda}_{l})$, we obtain a cloud of allowed points in $\mathbb{R}^{7}$. (We actually have allowed line segments between each pair of $\vec{\lambda}_{l}$ and $\vec{\lambda}_{u}$, but in practice, $|\vec{\lambda}_{l}-\vec{\lambda}_{u}|$ is tiny compared to the differences between $\vec{\lambda}$ at different points in dimension space, so the distinction is not important.) The bounds on $|\vec{\lambda}|^{2}$ come with extremal functionals containing interesting information about the spectrum of the 3d Ising CFT, which we will report on in future work.

### v2

@@ 1 Introduction
However, there are several reasons why we would like to go beyond these results. Firstly, it is not clear that the allowed region of Kos:2016ysd can be made systematically smaller by increasing the derivative order of bootstrap functionals. Relatedly, there are certain low-lying Regge trajectories of operators that are expected theoretically Fitzpatrick:2012yx ; Komargodski:2012ek ; Fitzpatrick:2015qma but not present in the extremal spectra studied in Simmons-Duffin:2016wlq . These include the $\sigma T]$, $\epsilon T]$, and $[TT]$ double-twist trajectories built out of the stress-energy tensor $T^{\mu\nu}$ (which can also be interpreted as triple-twist or quadruple-twist trajectories built out of $\sigma$). Furthermore, the spectra of Simmons-Duffin:2016wlq give no information about $\mathbb{Z}_{2}$-even operators of odd spin, parity-odd operators, or operator product expansion (OPE) coefficients involving multiple spinning operators (e.g. those appearing in $\langle TT\epsilon\rangle$ or $\langle TTT\rangle$).
@@
Table 1: Bounds on scaling dimensions and OPE coefficients in the 3d Ising CFT, using the numerical bootstrap for correlators of $\sigma$, $\epsilon$, and the stress tensor $T$. Here, $c_{T}/c_{B}$ is the coefficient of the stress-tensor two-point function (the “central charge”) normalized by the free boson value, and $n_{B}$ and $n_{F}$ are OPE coefficients in the stress-tensor three-point function. The OPE coefficient $\lambda_{TT\epsilon}$ is defined in section 2.1. Our bounds are computed using derivative order $\Lambda=43$. Bold-face errors are rigorous. We also include comparisons to previous bootstrap calculations and the best determinations we are aware of from other numerical methods. Note: the errors in Kos:2016ysd have been slightly underestimated, see footnote 2. ∗∗Note: the results for $(n_{B},n_{F})$ in Dymarsky:2017yzx assume a parity-odd scalar gap $\Delta_{-}\geq 10$. The weaker assumption $\Delta_{-}\geq 3$ gave $(n_{B},n_{F})=(0.919(18),0.027(18))$.
@@
Figure 1: Allowed regions in $(\Delta_{\sigma},\Delta_{\epsilon})$-space determined from the $T\sigma\epsilon$ system of correlators at $\Lambda=27,35,43$ (orange, green, and red), together with the allowed region determined from the $\sigma\epsilon$ system at $\Lambda=43$ (blue) Kos:2016ysd . [^2] The $\sigma\epsilon$ allowed region shown in this figure was re-computed for this work and is a more accurate version of the allowed region from Kos:2016ysd . Compared to Kos:2016ysd , we have benefitted from the OPE cutting surface algorithm Chester:2019ifh , improvements in computational efficiency, and more powerful computational resources. Our $\sigma\epsilon$ allowed region is slightly larger than the one in Kos:2016ysd but this does not significantly affect any of their results. We thank Slava Rychkov and Ning Su for discussions that motivated us to perform this calculation. On the left, we show a wider view where all the regions are visible. On the right, we zoom in on the $T\sigma\epsilon$ regions at $\Lambda=35$ and $\Lambda=43$. At $\Lambda=43$, the $T\sigma\epsilon$ allowed region is roughly $40$-$50\times$ smaller than the $\sigma\epsilon$ allowed region.
A natural continuation of these works is to consider a mixed-correlator system involving all nonvanishing 4-point functions of $\sigma$, $\epsilon$, and $T^{\mu\nu}$. While such a large system comes with technical challenges, they are not insurmountable and we will see that combining these constraints allows us to make precise predictions for CFT data involving the stress tensor, opening up a new era of precision in the numerical study of the 3d Ising CFT. Along the way we have significantly improved several software tools and algorithms used for the numerical bootstrap. For the impatient reader, a concise summary of our new results for scaling dimensions and OPE coefficients is provided in table 1. The corresponding bounds on the critical exponents $\eta,\nu$ are shown in table 2. We also provide a plot of the allowed island in $(\Delta_{\sigma},\Delta_{\epsilon})$-space (which is about 50 times smaller than the previous state of the art) in figure 1.
@@ 2.1 Three-point structures
where $V_{i}$ and $H_{ij}$ are standard embedding space structures Costa:2011mg , and $\widehat{T}$ is the stress-tensor with unit-normalized two-point function, defined in section 2.2 below. In the $\mathrm{SO}(3)_{r}$ basis, the structure $\texttt{ScalarParityEven}_{\Delta}$ is defined as
$\texttt{ScalarParityEven}_{\Delta}$ $=-\frac{(\Delta-5)(\Delta-3)}{4\sqrt{5}\Delta^{2}}|0,0\rangle+\frac{(5-\Delta)}{2\sqrt{14}\Delta}|2,2\rangle-\frac{(\Delta+2)}{4\sqrt{70}\Delta}|4,4\rangle,$ (3)
where $|j_{12},j_{123}\rangle$ are $\mathrm{SO}(3)_{r}$ basis elements. ($\texttt{ScalarParityEven}_{\Delta}$ is also given in a machine-readable format in the ancillary file.) The expressions for (2.1) under permutations of the 3 operators are given in (85).
@@
We emphasize that it is important to ensure that three-point structures for a given channel are linearly-independent for all allowed values of $\Delta$. [^4] We observed empirically that one only needs to ensure linear independence for the values of $\Delta$ allowed by unitarity and the gap assumptions. Presumably this happens because in SDPB the conformal blocks are eventually sampled at a large number of allowed values of $\Delta$ and polynomially interpolated. This interpolation will very accurately reproduce any degeneracies in the allowed region of $\Delta$, but will likely break any degeneracies outside of this region where the polynomial interpolation is of poor quality. Without care, it can happen that the space spanned by a naïve choice of conserved structures is degenerate at specific values of $\Delta$. This occurs, for example, if one solves the conservation equations in Mathematica in a generic basis.
@@ 2.2 Ward identities
$-\int_{\Sigma}dS_{\mu}x_{\nu}\langle\mathcal{O}\mathcal{O}T^{\mu\nu}\rangle=\Delta_{\mathcal{O}}\langle\mathcal{O}\mathcal{O}\rangle,$ (5)
@@ 2.3 Four-point structures and crossing equations
To describe four-point tensor structures we use the $q$-basis of Kravchuk:2016qvl , as it does not suffer from redundancies which plague embedding-space bases Costa:2011mg ; Dymarsky:2013wla . A $q$-basis four-point tensor structure is denoted by $q_{1}q_{2}q_{3}q_{4}]^{\pm}$, where $q_{i}\in\{-\ell_{i},-\ell_{i}+1,\dots,+\ell_{i}\}$ and $\ell_{i}$ is the spin of the operator at position $i$. The sign $\pm$ indicates the charge of the tensor structure under exchange of the cross-ratios $z\leftrightarrow\overline{z}$.
@@
$\Pi_{\langle TTTT\rangle}$ $=\{\text{id},(12)(34),(13)(24),(14)(23)\},$ (16)
@@
with analogous expressions for other orderings. Any $q$-basis structure can be symmetrized under $\Pi$, and we denote symmetrization of $q_{1}q_{2}q_{3}q_{4}]^{\pm}$ by $\langle q_{1}q_{2}q_{3}q_{4}\rangle^{\pm}$. Four-point functions can now be written as
@@ 3.2 Optimizing blocks_3d
In order to explain these improvements, we must give a brief overview of the function of blocks_3d. A single run of blocks_3d computes multiple conformal blocks, which all share the same external operator representations, four-point structure label, as well as the $j_{12}$ and $j_{43}$ parts of the three-point function labels. The program outputs conformal blocks with all possible values of $j_{120}$ and $j_{430}$, where the index 0 refers to the exchanged operator, as well as all values of the exchanged spin $j$ up to a cutoff $j_{\text{max}}$.
@@
[…]
@@ 4.1 Review: cutting surface algorithm
Importantly, this constraint is unchanged if we rescale $\vec{\lambda}$ by a non-zero real number. Therefore, it only depends on the equivalence class $\vec{\lambda}]$ of $\vec{\lambda}$ in $\mathbb{R}\mathbb{P}^{4}$ and we should treat $\vec{\lambda}]$ as the new parameter. If we are able to exclude all possible values of $\vec{\lambda}]$ for a fixed $(\Delta_{\sigma},\Delta_{\epsilon})$, then the point $(\Delta_{\sigma},\Delta_{\epsilon})$ is excluded in the dimension space. As shown in Kos:2016ysd , this procedure yields a significantly smaller allowed region for $(\Delta_{\sigma},\Delta_{\epsilon})$ than using (31).
How can one exclude all possible values of $\vec{\lambda}]$ in a finite amount of time? The approach used in Kos:2016ysd was to discretize $\mathbb{R}\mathbb{P}^{1}$ (their $\vec{\lambda}$ had only two components) using a finite grid. If all points in the grid are excluded, one can be reasonably confident that there are no allowed points in $\mathbb{R}\mathbb{P}^{1}$ if the grid is fine enough. This approach worked well in Kos:2013tga because their parameter space $\mathbb{R}\mathbb{P}^{1}$ was one-dimensional. In our case the parameter space is $\mathbb{R}\mathbb{P}^{4}$ and discretization appears computationally infeasible.
This problem was solved in Chester:2019ifh with the “cutting surface” algorithm. The key observation is simply that a finite set of functionals $\alpha_{I}^{(k)}$, $k=\{1,2,\dots,N\}$, can be enough to exclude the full parameter space $\mathbb{R}\mathbb{P}^{4}$. Indeed, existence of a functional $\alpha_{I}^{(k)}$ implies that all $\vec{\lambda}]$ which satisfy
@@
If (36) holds, then all $\vec{\lambda}]$ are excluded. If (36) doesn’t hold, then we can take a $\vec{\lambda}_{N+1}]$ that is not in $\bigcup_{k=1}^{N}E_{k}$, i.e.
@@
If such a functional $\alpha_{I}^{(N+1)}$ is found, we can add it to the list, increment $N$, and check (36) again. If a functional cannot be found, it means that $\vec{\lambda}_{N+1}]$ is allowed and so are the current values of $(\Delta_{\sigma},\Delta_{\epsilon})$. Typically, this process converges in a number of steps which is much smaller than the number of points in a fine enough grid on $\mathbb{R}\mathbb{P}^{4}$. It also has the benefit of not relying on a discretization.
The technical content of the cutting surface algorithm Chester:2019ifh is in the procedure for finding a solution $\vec{\lambda}_{N+1}]$ to (37) or determining that no solution exists. [^17] When several solutions to (37) can be found, the cutting surface algorithm will try to select a solution that is the furthest away from the sets $E_{k}$ in order to speed up the convergence of the whole process. Abstractly, this can be formulated as the following problem: given quadratic forms $Q_{0},Q_{1},\dots,Q_{N}$, determine whether there exists a solution $\vec{\lambda}\in\mathbb{R}^{n+1}$ to the equations
@@
the quadratic form $Q_{0}$ is specified manually to restrict the search to an ellipse in $\mathbb{R}\mathbb{P}^{n+1}$. This bounding ellipse is chosen conservatively to include the region where the allowed points $\vec{\lambda}]$ are expected to be. See section 5.2 for details. In essence, one of our assumptions is that if any $\vec{\lambda}]$ is allowed at all, it is deep inside the region defined by $\vec{\lambda}^{T}Q_{0}\vec{\lambda}<0$. The technical meaning of this will be clarified below.
@@ 4.2 Certification algorithm
Out of the 38 disallowed points in $(\Delta_{\sigma},\Delta_{\epsilon})$ space that we computed at $\Lambda=43$, 30 points were certified by $\mathcal{B}[\delta c^{1}]=\emptyset$, 1 point by $\mathcal{B}[\delta c^{2}]=\emptyset$, and 2 points by $\mathcal{B}[\delta c^{3}]=\emptyset$. This produces certificates for 33 out of 38 points.
@@ 5.1.1 Bounds on scalar gaps
Figure 4: Bound on the allowed gaps for parity-even and -odd scalars in the single-correlator $\langle TTTT\rangle$ system. Successively interior regions correspond to higher derivative order $\Lambda$. We can see that by $\Lambda=43$ the bounds are fairly well converged and therefore are hard to distinguish visually at this scale. The known values for these gaps in the free scalar (FS) theory, Majorana free fermion ($F$) theory, stress tensor mean-field theory (MFT), $O(N)$ CFT in the large-NN limit ($O(\infty)$), and Ising CFT are indicated.
@@ 5.2 Mixed correlators of $T,\sigma,\epsilon$
++ ++ 0 $\Delta\geq 3$ $\Delta=3.82951(61)$ Reehorst:2021hmp
- ++ 0 $\Delta\geq 3$ $\Delta=5.262(89)$ Reehorst:2021hmp
++ - 0 $\Delta\geq 3$ $\Delta\gtrsim 10$ Zhu:2022gjc
++ ++ 2 $\Delta\geq 4$ $\Delta=5.499(17)$ Reehorst:2021hmp
@@
Table 6: Our gap assumptions for operators other than $\sigma,\epsilon,T$, used for computing the allowed island in the 3d Ising CFT. Essentially, we assume that all scalars (other than $\sigma$, $\epsilon$) are irrelevant, and that there is a gap of 11 above the stress tensor. We also assume a small twist gap in all other sectors ($\tau=\Delta$-$\ell$). In the right-hand column, we display the expected actual values for these gaps in the 3d Ising CFT. In the last row of the right-hand column, the twist gap comes from the $\mathbb{Z}_{2}$-even, parity-even, spin-4 sector, which has $\Delta_{C_{\mu\nu\rho\sigma}}=5.022665(28)$. Clearly our gap assumptions are very conservative.
[…]
@@
To compute bounds at higher derivative order, we followed a similar procedure to the computations of bootstrap islands in the $O(2)$ model Chester:2019ifh , $O(3)$ model Chester:2020iyt , and GNY models Erramilli:2022kgp . For each point in $(\Delta_{\sigma},\Delta_{\epsilon})$-space, we use the OPE “cutting surface” algorithm of Chester:2019ifh to find an allowed direction for the vector $\vec{\lambda}$. The output of this algorithm is either an equivalence class $\vec{\lambda}]\in\mathbb{RP}^{4}$ for which the corresponding SDP is primal feasible, or evidence that no such $\vec{\lambda}]$ exists (see section 4 for details). In the former case, we say that the point $(\Delta_{\sigma},\Delta_{\epsilon})$ is “allowed,” and in the latter case we say it is “disallowed.”
@@
$\Delta_{\sigma}$ $=0.5181488106917789+8.496193518377354\cdot 10^{-8}x-5.774118868365222\cdot 10^{-9}y,$
$\Delta_{\epsilon}$ $=1.412625302544141+1.061340020903086\cdot 10^{-6}x+4.622272819035116\cdot 10^{-10}y.$ (53)
@@
The OPE cutting surface algorithm requires an initial bounding ellipse to work efficiently. We determine it by working our way up in derivative order $\Lambda$, with a method similar to our strategy for finding affine coordinates $(x,y)$. At a given value of $\Lambda$, we collect together all the allowed points in OPE space $\vec{\lambda}]$ and use principal component analysis (in an affine patch of $\mathbb{RP}^{4}$) to find an ellipse that contains them all. Then we expand this ellipse by an $O(1)$ factor, say $4$. Finally, we use the expanded ellipse as the bounding ellipse for the next higher value $\Lambda^{\prime}>\Lambda$. As a consistency check, we observe afterwards that the allowed OPE points $\vec{\lambda}]$ at the higher $\Lambda^{\prime}$ cluster in a small region deep in the interior of the bounding ellipse, indicating that the results are independent of the choice of bounding ellipse. As an example, the bounding ellipse at $\Lambda=43$ is defined by $\vec{\lambda}^{T}Q_{0}\vec{\lambda}\leq 0$, where $Q_{0}$ is given by the matrix
@@
In figure 8, we show a plot of the allowed points $\vec{\lambda}]$ at $\Lambda=43$, compared to this bounding ellipse (which is a sphere of radius $4$ in the coordinates of the plot). As claimed, all the allowed points are deep in the interior.
Figure 8: Allowed values of $\vec{\lambda}]$ found at $\Lambda=43$, compared to the initial bounding ellipse, in coordinates where the bounding ellipse is a sphere of radius 44. Here, we show just 3 out of 4 coordinates of $\vec{\lambda}]$ — the other coordinate is similar. All the allowed points are deep inside the bounding ellipse.
As mentioned above, each allowed point in $(\Delta_{\sigma},\Delta_{\epsilon})$-space comes with an equivalence class $\vec{\lambda}]\in\mathbb{RP}^{4}$, determined by the OPE cutting surface algorithm. Next, for each allowed point, we compute upper and lower bounds on the norm-squared $|\vec{\lambda}|^{2}$ using the standard procedure for bounding OPE coefficients Caracciolo:2009bx . This results in pairs of OPE coefficient vectors $\vec{\lambda}_{u},\vec{\lambda}_{l}$ with norms given by the upper/lower bounds on $|\vec{\lambda}|$. Collecting together the data $(\Delta_{\sigma},\Delta_{\epsilon},\vec{\lambda}_{u})$ and $(\Delta_{\sigma},\Delta_{\epsilon},\vec{\lambda}_{l})$, we obtain a cloud of allowed points in $\mathbb{R}^{7}$. (We actually have allowed line segments between each pair of $\vec{\lambda}_{l}$ and $\vec{\lambda}_{u}$, but in practice, $|\vec{\lambda}_{l}-\vec{\lambda}_{u}|$ is tiny compared to the differences between $\vec{\lambda}$ at different points in dimension space, so the distinction is not important.) The bounds on $|\vec{\lambda}|^{2}$ come with extremal functionals containing interesting information about the spectrum of the 3d Ising CFT, which we will report on in future work.